Home | New Request

Facility Management Requests Status Page

SEARCH BY

Request
ID
Date
Submitted
Submitted
By
Lab/Group Description Status Comment
Posted
Comment
55336 11/08/16 Granlund, Eric Granlund, Eric  1946261 - B4 Lat - 2 lights outside in hallway Reported to Campus  
55055 10/03/16 Becknell, Nigel Yang, Peidong  1940319 - 315/A Hild - 19 Lights out Reported to Campus  
54820 08/22/16 Brateng, Michael   1930469 - 610 Lat - 10/30 lights out Reported to Campus  
54799 08/17/16 Hung, Kevin   Door handle of 917 broken Reported to Campus  
54631 07/22/16 Sutton, Suzanne   1924992 - 101 Lew - Weatherstrip windows Reported to Campus  
56248 04/21/17 Chan, Leslie Maboudian, Roya  B78 Tan - Ceiling leaking Reported to Campus  
56239 04/19/17 Lee, Taegyo Hartwig, John  702 Lat - F12000 has 50 fpm In Progress  
56227 04/18/17 Smith, Matthew Francis, Matthew  3158653 - 743 Lat - balcony flooded, not draining. Drain clogged. Reported to Campus  
56218 04/17/17 Hardy, Melissa Sarpong, Richmond  3158744 - 842 Lat - Ceiling leak left of FH F12459 Reported to Campus  
56211 04/15/17 Hu, Xirui Maimone, Thomas  3156396 - 916 Lat - Eye wash in lab does not work Reported to Campus  
56207 04/14/17 Swabeck, Joseph Alivisatos, Armand  3155083 - D79 Hild - Fire extinguishers out of date Reported to Campus  
56192 04/13/17 Kerschgens, Isabel Sarpong, Richmond  3160922 - 842 Lat - F12459 goes into alarm due to low fpm In Progress  
56173 04/10/17 Gibson, Natalie Leone, Stephen  3152472 - D44 Hildebrand - 5 lights out Reported to Campus  
56157 04/05/17 Jang, Hee-Jong Beeler, Rose  3149001 - 2nd fl Lat - Wrr has 1 light out Reported to Campus  
56147 04/03/17 Winslow, Charles Toste, F. Dean  3148378 - 605 Lat - F12247 making loud drilling sound from above Reported to Campus  
56144 04/03/17 Leung, Samuel Iglesia, Enrique  3148452 - 275 Tan - 6 lights out Reported to Campus  
56123 03/28/17 Concepcion, Agnes Concepcion, Agnes  3141147 - 100A Hild - 2 lights out Reported to Campus  
56106 03/27/17 Kerschgens, Isabel Sarpong, Richmond  3138458 - 842 Lat - F12459 is making loud noise from top Reported to Campus  
56093 03/24/17 Bohan, Patrick Toste, F. Dean  3137323 - 605 Lat - balcony door won't open Reported to Campus  
56062 03/20/17 Gardner, David Maboudian, Roya  3127041 - B52 Tan - The black flat that is underneath the door to is coming loose. Reported to Campus  
56055 03/16/17 Winslow, Charles Toste, F. Dean  3125533 - 605 Lat - House Vac Knob Broken on Hood Reported to Campus  
56044 03/13/17 Hong, Cynthia Toste, F. Dean  3123349 - 505 Lat - F11964 sash needs greasing Reported to Campus  
56001 03/07/17 Sareshwala, Anjum   3119175 - 330 Lat - Too cold In Progress  
55989 03/06/17 Becknell, Nigel   3118838 - 330 Hild - 6/12 lights out Reported to Campus  
55988 03/06/17 Becknell, Nigel   3118839 - 334 Hild - 9/12 lights out Reported to Campus  
55997 03/06/17 Boreen, Michael Arnold, John  3119170 - 510A Lat - 3 lights out Reported to Campus  
55992 03/06/17 Shusterman, Alexis Cohen, Ronald  3119171 - B76 Hild - 10/16 lights out Reported to Campus  
55974 03/02/17 Hartman, Emily Francis, Matthew  1934282 - 905/903 Lat - 13/30 lights out Reported to Campus  
55914 02/24/17 Wilson, Emery   3109157 - 772 Tan - Night security light out in hall Reported to Campus  
55907 02/23/17 Kapelewski, Matthew Long, Jeffrey  B30 Hild - Ventilation not working In Progress  
55889 02/21/17 Madsen, Catherine   3101615 - 791 Tan - 1 ceiling light out Reported to Campus  
55822 02/08/17 Massen, Inna   3062763 - 810 Lat - Roof leak Reported to Campus  
55801 02/04/17 Han, Seojung Toste, F. Dean  3088164 - 619 Lat - 4 lights out Reported to Campus  
55771 02/01/17 Gardner, David Maboudian, Roya  3079620 - B72 Tan - 1 light out above desk 5 Reported to Campus  
55767 01/31/17 Armstrong, Laura Baranger, Anne  3079614 - 318 Lew - Windows not closing/sealing all the way Reported to Campus  
55738 01/30/17 Boering, Kristie   3079401 - BG3 - 6 lights out Reported to Campus  
55700 01/25/17 Jiang, Henry Long, Jeffrey  3078883 - B54A Hild - 1 light out Reported to Campus  
55660 01/20/17 Monahan, Daniele Fleming, Graham  3062819 - B75 Hild - 1 light flickering Reported to Campus  
55661 01/20/17 Kulkarni, Rishikesh Miller, Evan  3069644 - 407 Hild - Leak between roof and duct In Progress